OpenStudy (anonymous):

write a linear factorization of x^3+4x-5

OpenStudy (muzzack):

Possible roots = +/-1, +/-2, +/-5 Let's test if 2 is a possible zero. 8 - 24 + 26 - 10 -2 + 2 = 0 2 is one of the zeroes. Use synthetic division and divide the expression by (x-2). Then multiply that by (x-2). (x^3 - 6x^2 + 13x - 10) / (x-2) = x^2 - 4x + 5 (x-2)(x^2-4x+5) Use quadratic formula to find the zeros found in \(x^2-4x+5\) x = 2+/-i Zeroes: 2, 2+/-i

OpenStudy (anonymous):

But two doesn't work for that

OpenStudy (anonymous):

The only real zero should be 1